OpenStudy (anonymous):

I'm having a hard time understanding the similarity between function notation and evaluating a polynomial.

OpenStudy (anonymous):

so wats ur question? plz state it clearly so there is no confusion as to wat your asking

OpenStudy (anonymous):

What are the similarities and differences between evaluating a function notation and evaluating a polynominal

OpenStudy (anonymous):

by example: if \[f(x)=x^2-2x+3\] then \[f(3)=3^2-2\times 3 + 3\]

OpenStudy (anonymous):

so evaluating a polynomial function at a number is the same as "evaluating the polynomial". on the other hand the phrase "evaluating a function notation" has no meaning that i can figure out

OpenStudy (anonymous):

K thanks